OpenStudy (anonymous):

The width w of a rectangular swimming pool is 7x2. The area A of the pool is 7x3 – 42x2. What is an expression for the length of the pool? Show your work.

Mehek (mehek14):

\(\dfrac{7x^3-42x^2}{7x^2}\)

Mehek (mehek14):

do you know how to divide this?

OpenStudy (anonymous):

Do i divide the the numerators with the same denominator? If so i think i can

OpenStudy (anonymous):

because 7 can only be divided by 7 once

Mehek (mehek14):

yes like this \(\dfrac{7x^3}{7x^2}\)

Mehek (mehek14):

correct so 7/7 = 1 how about \(\dfrac{x^3}{x^2}\)

OpenStudy (anonymous):

Think it's just x

Mehek (mehek14):

correct

Mehek (mehek14):

now \(\dfrac{42x^2}{7x^2}\)

OpenStudy (anonymous):

Is it 6, 6+x = 6x @Mehek14

Mehek (mehek14):

so the length is \(x+6\)

OpenStudy (anonymous):

Okay had it written wrong Thanks Mehek14

Mehek (mehek14):

\((7x^2)(x+6) = 7x^3 + 42x^2\)

Mehek (mehek14):

wait it's -6

OpenStudy (anonymous):

Oh? how did it get negative?

Mehek (mehek14):

\(7x^2*(x-6)=7x^3-42x^2\)

Mehek (mehek14):

because -42/7 = -6

OpenStudy (anonymous):

Oh, didn't notice the 42 was negative i would of been way off

Mehek (mehek14):

me neither :P